the pressure at point A is 25 psi. All fluids are at 20 C. What is the

air pressure in the closed chamber B, in Pa? SAE 30 oil ρ= 891 kg/m3


1
Expert's answer
2021-03-22T08:37:17-0400


Take γ = 9790 N/m3 for water, 8720 N/m3 for SAE 30 oil, and (1.45)(9790) = 14196 N/m3 for the third fluid. Convert the pressure at A from 25 lbf/in2 to 172400 Pa.

Compute hydrostatically from point A to point B:

pA + "\sum\gamma" h = 172400 Pa - (9790 N/m3)(0.04 m) + (8720 N/m3)(0.06 m) - (14196 N/m3)(0.10 m) =

= pB = 71100 Pa / 47.88 / 144 = 24.8 psi